OpenStudy (anonymous):

What is the slope and y-intercept of the graph of 3x + 6y = 36?

OpenStudy (anonymous):

slope -1/2 and y intercept 1/6

OpenStudy (shaik0124):

6y=-3x+36 y=-3/6x+36/6 y=-1/2x+6 as u know that y=mx+c where m is slope so m=-1/2 on comparing both equations

OpenStudy (shaik0124):

@rosecastillio understood

OpenStudy (anonymous):

and the y intercept is (0,6) correct? @shaik0124

OpenStudy (shaik0124):

slope intercept form: y = mx+b slope = m = rise/run y-intercept = b y intercept is 6

OpenStudy (shaik0124):

@rosecastillio got it